OpenStudy (anonymous):

Theorem: there are irrational a and b such that a^b is rational. Is the following proof correct? (note: use the fact that √(2) is irrational) Proof: either (√(2))^√(2) is rational or irrational. Case 1: suppose (√(2))^√(2) is rational. Let a = b = √(2), a^b = (√(2))^√(2) is rational by assumption Case 2: suppose (√(2))^√(2) is irrational. Let a = (√(2))^√(2), and b = √(2). Then a^b = [(√(2))^√(2)]^√(2) = [√(2)]^2 = 2, which is rational.

OpenStudy (anonymous):

I think this is like a circular reason. In second case, (√(2))^√(2) is assumed to be rational but then the conluclusion says it's rational. In the first case it's assumed (√(2))^√(2) is rational, but this is the conclusion of the theorem.

OpenStudy (anonymous):

It's not circular reasoning.

OpenStudy (anonymous):

The second case assumes (√(2))^√(2) is irrational.

OpenStudy (anonymous):

i think i was trying to prove sqrt(2)^sqrt(2) is rational, which is not what the theorem says :D

OpenStudy (anonymous):

The theorem says \(\exists a,b \in \mathbb R \setminus \mathbb Q \quad a^b \in \mathbb Q\). This is much different than \(\forall a,b \in \mathbb R \setminus \mathbb Q \quad a^b \in \mathbb Q\).